OpenStudy (anonymous):

Factor for solutions: X^3 -64=0 Is it X=4. I'm just making sure if it is?

OpenStudy (anonymous):

yes!! X^3 -64=0 x^3 = 64 x^3 = 4 x 4 x 4 = 64 x = 4

OpenStudy (mathmath333):

here u will get 3 solutions including x=4 and two imaginary solutions

OpenStudy (mathmath333):

\(\large\tt \color{black}{x^3-4^3=0}\) apply the formula \(\large\tt \color{black}{a^3-b^3=(a-b)(a^2+ab+b^2)}\)

OpenStudy (anonymous):

Would I get 3 solutions for 2x^2+10x+8 or just 2 solutions

OpenStudy (anonymous):

for that i got (X+1) (x+4)

OpenStudy (mathmath333):

for this 2x^2+10x+8 it will be two as the power of x is 2

OpenStudy (anonymous):

would it be (X+1) (x+4)

OpenStudy (mathmath333):

\(\huge\tt \color{black}{= 2(x+1) (x+4)}\)

OpenStudy (anonymous):

Oh okay

OpenStudy (anonymous):

Now I have 1 more. For this problem, X^3+5X^2+6X=0. Would it be 3 solutions and if so would it be X=0, -2, -3

OpenStudy (mathmath333):

u can check it urself by putting each solution in the equation if it is equal to zero then u r correct

OpenStudy (mathmath333):

the power of x is three so here will be three solutions

OpenStudy (anonymous):

[k great thanks

OpenStudy (anonymous):

Sorry but for X^3 -64=0, Would the solutions be X= 4, 2-2i√ 3, 2+2i√ 3

OpenStudy (mathmath333):

yes of cos it would be